# Gradient Extrapolation for Debiased Representation Learning

### v1

@@ Appendix C GERNE versus an equivalent sampling and weighting approach
    A binary classification task where the number of classes equals the number of attributes (i.e. $K=A=2$).
@@
    The attributes are known, and the classes are balanced. (i.e. $|\mathcal{X}_{y=1}|=|\mathcal{X}_{y=2}|$).
@@
    The majority of samples which hold the spurious correlation in each class are aligned with the class label (i.e. $|\mathcal{X}_{y,a=y}|>|\mathcal{X}_{y,a\neq y}|$).
@@
    The dataset is highly biased. In other words, $\frac{|\mathcal{X}_{y,a\neq y}|}{|\mathcal{X}_{y,a=y}|}\ll 1$.
  5. 5.
    In a highly biased dataset, best performance is coupled with overpresenting the minority (conflicting samples according to assumption 1.) in early stages of training. Therefore, an overfitting on the minority is expected before the overfitting on the majority.
We refer to the expected loss of the majority (aligned) samples as $\mathcal{L}_{A}$, and the expected loss of the minority (conflicting) as $\mathcal{L}_{C}$. For GERNE, we sample two batches: biased and less biased batch, each of size $B$. From Eq. 5, $\mathcal{L}_{ext}$ can be written as:
$\mathcal{L}_{ext}=(1+\beta)\cdot\mathcal{L}_{lb}-\beta\cdot\mathcal{L}_{b}$ (17)
Since the biased batch reflects the inherent bias present in the dataset, under the third assumption, we can approximate $\mathcal{L}_{b}$ by $\mathcal{L}_{A}$, neglecting the loss on the very few conflicting samples in the batch. Therefore, we have:
$\mathcal{L}_{b}\approx\mathcal{L}_{A}$ (18)
Following the third assumption and the conditional attribute distribution in Eq. 4, we can approximate the composition of the less biased batch as follows: a proportion of $(1-\frac{c}{2})$ of the samples in the less biased batch are drawn from the aligned samples, while a proportion of $\frac{c}{2}$ of the samples from the minority group. this leads to the following approximation:
$\mathcal{L}_{lb}\approx(1-\frac{c}{2})\cdot\mathcal{L}_{A}+\frac{c}{2}\cdot\mathcal{L}_{C}$ (19)
@@
We consider the following ’SW’ approach:
  • Sampling step : we sample an ’SW’ batch of size $B$ similar to the less biased batch in GERNE. Where $(1-\frac{c}{2})$ of the batch samples are from the majority group (aligned samples) and $\frac{c}{2}$ from minority (conflicting samples).
@@
    $\mathcal{L}_{sw}=w\cdot\mathcal{L}_{A}+(1-w)\cdot\mathcal{L}_{C},w=\frac{2-c\cdot(1+\beta)}{2}$ (21)
    where $\mathcal{L}_{A}$ is computed over aligned samples in the ’SW’ batch and $\mathcal{L}_{C}$ is computed over the conflicting samples.
Let’s compute the variance of the two losses:
@@
$\quad+2\cdot w\cdot(1-w)\cdot Cov(\mathcal{L}_{A}^{1-c/2},\mathcal{L}_{C}^{c/2})$ (22)
where $Var(\mathcal{L}^{m})$ means the variance computed over $m\cdot B$ samples where $B$ is the batch size. For simplicity, we refer to $Var(\mathcal{L}^{1})\text{ as }Var(\mathcal{L})$
Following the fourth assumption, when the model overfits on the conflicting samples (i.e. $\mathcal{L}_{C}\approx 0$), we can approximate both $Var(\mathcal{L}_{C}),Cov(\mathcal{L}_{A},\mathcal{L}_{C})$ to zero. Therefore:
@@
$\quad(\frac{2-c\cdot(1+\beta)}{2})^{2}\cdot\frac{2}{2-c}\cdot Var(\mathcal{L}_{A})$ (23)
@@
$\quad\geq((1+\beta)\cdot\sqrt{Var(\mathcal{L}_{lb})}-\beta\cdot\sqrt{Var(\mathcal{L}_{b})})^{2}$ (24)
@@
Note that the the inequality reduces to an equality in Appendix C if $Cov(\mathcal{L}_{lb},\mathcal{L}_{b})=\sqrt{Var(\mathcal{L}_{lb})}\cdot\sqrt{Var(\mathcal{L}_{b})}$.
@@
The covariance term ($Cov$) can be controlled by the number of shared samples between the biased and less biased batches. If all the aligned samples in the less biased batch are included in the sampled biased batch (i.e. the less biased batch is created by replacing some aligned samples by conflicting ones), we get maximum value for the $Cov$.
From Eq. 19:
$Var(\mathcal{L}_{lb})\approx(1-\frac{c}{2})^{2}\cdot Var(\mathcal{L}_{A}^{1-c/2})=(1-\frac{c}{2})\cdot Var(\mathcal{L}_{A})$ (25)
And from Eq. 18
$Var(\mathcal{L}_{b})\approx Var(\mathcal{L}_{A})$ (26)
Finally, substituting Eq. 25 and Eq. 26 in Appendix C:
$Var(\mathcal{L}_{ext})\geq((1+\beta)\cdot\sqrt{1-\frac{c}{2}}-\beta)^{2}\cdot Var(\mathcal{L}_{A})$ (27)
@@
According to the fourth assumption, we are interested in the range where $c\cdot(\beta+1)\geq 1$. Using the limits of $\beta$ defined in Eq. 9, we obtain $\beta\in\left[\frac{1-c}{c},\frac{2-c}{c}\right]$. As $\beta\to\frac{2-c}{c}$, the representation of aligned samples simulates a vanishing representation (according to Eq. 8) in the sampled batches, which leads to $\mathcal{L}_{A}>0$. Assuming a limited and non-vanishing variance $Var(\mathcal{L}_{A})$ (i.e. the model outputs a non-constant prediction for samples from the majority group), we have:
@@
$\beta\to\frac{2-c}{c}\implies Var(\mathcal{L}_{sw})\approx 0,Var(\mathcal{L}_{ext})\neq 0$ for $c\in(0,1]$. This non-vanishing variance of GERNE’s loss, if controlled with tuning $\beta$ to ensure stability, gives the model the chance of escape sharp minima similar to gradient random perturbation [1] and therefore, improve generalization [33, 23, 34].
Appendix D Algorithm 2.
@@
Input: $\mathcal{X}_{y}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$, $f$ with initial $\theta=\theta_{0},\tilde{\theta}=\tilde{\theta_{0}}$ (parameters of the biased model $\tilde{f}$), ##\## epochs $E$, batch size per class label $B$, ##\## classes $K$, ##\## attributes $\tilde{A}=2$, learning rate $\eta$.
1:Training $\tilde{f}$ on biased batches with class balanced accuracy CBA $=\frac{1}{K}\sum_{y\in\mathcal{Y}}\mathbb{P}_{x|y}(y=\arg\max_{y^{\prime}\in\mathcal{Y}}\tilde{f}_{y^{\prime}}(x))\quad$ as the evaluation metric for model selection.
2:Select a threshold $t$ and create the pseudo-groups $\tilde{\mathcal{G}}$ by following the steps in Sec. 4.2.
3:Follow Algorithm 1. with: $\mathcal{G}\leftarrow\tilde{\mathcal{G}}$.
@@
Creating both biased and less biased batches using the pseudo-groups $\mathcal{\tilde{G}}$, and with $\beta$ as hyperparameter, we can simulate batches with a more controllable conditional attribute distribution. Specifically, for $(y,a)\in\mathcal{G}$, we can achieve scenarios where $p_{ext}(a|y)>\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\text{ or }p_{ext}(a|y)<\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$ as opposed to Eq. 13.
Proof. We define $\tilde{\alpha}_{y\tilde{a}}$ the same way as in Eq. 2 for the created pseudo-groups: $\tilde{\alpha}_{y\tilde{a}}=\frac{|\mathcal{X}_{y,\tilde{a}}|}{|\mathcal{X}_{y}|}$. For a constant $c$, we create the less biased batch as in Eq. 4:
$p_{lb}(\tilde{a}|y)=\tilde{\alpha}_{y\tilde{a}}+c\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}}).$ (28)
@@
$p_{ext}(\tilde{a}|y)=\tilde{\alpha}_{y\tilde{a}}+c\cdot(\beta+1)\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}}).$ (29)
@@
$p_{ext}(a|y)=\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\tilde{\alpha}_{y\tilde{a}}\cdot p(a|\tilde{a},y)+c\cdot(\beta+1)\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}})\cdot p(a|\tilde{a},y).$ (31)
For $p(a|\tilde{a}=1,y)\neq p(a|\tilde{a}=2,y)$ and $\tilde{\alpha}_{y1}\neq\frac{1}{2}$, to make $p_{ext}(a|y)=p$ for some $p\in[0,1]$, we can choose:
$\beta=\frac{p-\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\tilde{\alpha}_{y\tilde{a}}\cdotp(a|\tilde{a},y)}{\sum_{\tilde{a}\in\tilde{\mathcal{A}}}c\cdot(\frac{1}{2}-\tilde{\alpha}_{y\tilde{a}})\cdotp(a|\tilde{a},y)}-1.\blacksquare$ (32)
Discussion. When $\tilde{\alpha}_{y1}=\frac{1}{2}$, our algorithm is equivalent to sampling uniformly from $\mathcal{X}_{y}$ and equally from classes. When $p(a|\tilde{a}=1,y)=p(a|\tilde{a}=2,y)$, it implies that $\tilde{f}$ has distributed the samples with attribute $a$ and class $y$ equally between the two pseudo-groups. However, in practice, this is precisely the scenario that $\tilde{f}$ is designed to avoid. Specifically, if $a$ represents the presence of spurious attributes (i.e., the majority group), it is likely that $p(a|\tilde{a}=1,y)<p(a|\tilde{a}=2,y)$. Conversely, when $a$ represents the absence of spurious features (i.e., the minority group), we would expect $p(a|\tilde{a}=1,y)>p(a|\tilde{a}=2,y)$. In fact, $\tilde{f}$ is explicitly trained to exhibit a degree of bias, which inherently disrupts the equality above.

### v2

@@ Appendix B GERNE Versus an Equivalent Sampling and Weighting Approach
    A binary classification task where the number of classes equals the number of attributes (i.e., $K=A=2$).
@@
    The attributes are known, and the classes are balanced (i.e., $|\mathcal{X}_{y=1}|=|\mathcal{X}_{y=2}|$).
@@
    The majority of samples that hold the spurious correlation in each class are aligned with the class label (i.e., $|\mathcal{X}_{y,a=y}|>|\mathcal{X}_{y,a\neq y}|$), and the dataset is highly biased (i.e., $\frac{|\mathcal{X}_{y,a\neq y}|}{|\mathcal{X}_{y,a=y}|}\ll 1$).
@@
    In a highly biased dataset, best performance is coupled with overrepresenting the minority groups (as illustrated in Fig. 2) in early stages of training. Therefore, an overfitting on the minority groups is expected before the overfitting on the majority groups.
We refer to the expected loss of the majority samples as $\mathcal{L}_{A}$, and the expected loss of the minority as $\mathcal{L}_{C}$. For GERNE, we sample two batches: biased and less biased batch, each of size $B$. From Eq. 5, $\mathcal{L}_{ext}$ can be written as:
$\mathcal{L}_{ext}=(1+\beta)\cdot\mathcal{L}_{lb}-\beta\cdot\mathcal{L}_{b}.$ (16)
Since the biased batch reflects the inherent bias present in the dataset, under the third assumption, we can approximate $\mathcal{L}_{b}$ by $\mathcal{L}_{A}$, neglecting the loss on the very few samples of the minority group in the batch. Therefore, we have:
$\mathcal{L}_{b}\approx\mathcal{L}_{A}.$ (17)
Following the third assumption and the conditional attribute distribution in Eq. 4, we can approximate the composition of the less biased batch as follows: a proportion of $(1-\frac{c}{2})$ of the samples in the less biased batch are drawn from the aligned samples (majority group), while a proportion of $\frac{c}{2}$ of the samples from the minority group. This leads to the following approximation:
$\mathcal{L}_{lb}\approx(1-\frac{c}{2})\cdot\mathcal{L}_{A}+\frac{c}{2}\cdot\mathcal{L}_{C}.$ (18)
@@
We consider the following “SW” approach:
  • Sampling step: we sample an “SW” batch of size $B$, analogous to the less biased batch in GERNE, where $(1-\frac{c}{2})$ of the samples are from the majority group (aligned samples) and $\frac{c}{2}$ from minority group.
@@
    $\mathcal{L}_{sw}=w\cdot\mathcal{L}_{A}+(1-w)\cdot\mathcal{L}_{C},w=\frac{2-c\cdot(1+\beta)}{2},$ (20)
    where $\mathcal{L}_{A}$ is computed over the samples of the majority group in the “SW” batch and $\mathcal{L}_{C}$ is computed over the samples of the minority group.
Both $\mathcal{L}_{ext}$ and $\mathcal{L}_{sw}$ in Eq. 19, Eq. 20 are equivalent in expectation. Let’s compute the variance of the two losses:
@@
$\quad+2\cdot w\cdot(1-w)\cdot Cov(\mathcal{L}_{A}^{1-c/2},\mathcal{L}_{C}^{c/2}),$ (21)
[…]
@@
$\quad(\frac{2-c\cdot(1+\beta)}{2})^{2}\cdot\frac{2}{2-c}\cdot Var(\mathcal{L}_{A}).$ (22)
@@
$\quad\geq((1+\beta)\cdot\sqrt{Var(\mathcal{L}_{lb})}-\beta\cdot\sqrt{Var(\mathcal{L}_{b})})^{2}.$ (23)
@@
Note that the inequality reduces to an equality in Appendix B if $Cov(\mathcal{L}_{lb},\mathcal{L}_{b})=\sqrt{Var(\mathcal{L}_{lb})}\cdot\sqrt{Var(\mathcal{L}_{b})}$.
@@
The covariance term $Cov(.,.)$ can be controlled by the number of shared samples between the biased and less biased batches. If all the aligned samples in the less biased batch are included in the biased batch (i.e., the less biased batch is created by replacing some samples of the majority group with samples from the minority group), we then maximize $Cov(.,.)$. From Eq. 18, we can write:
$Var(\mathcal{L}_{lb})\approx(1-\frac{c}{2})^{2}\cdot Var(\mathcal{L}_{A}^{1-c/2})=(1-\frac{c}{2})\cdot Var(\mathcal{L}_{A}),$ (24)
and from Eq. 17:
$Var(\mathcal{L}_{b})\approx Var(\mathcal{L}_{A}).$ (25)
Finally, substituting Eq. 24 and Eq. 25 in Appendix B:
$Var(\mathcal{L}_{ext})\geq((1+\beta)\cdot\sqrt{1-\frac{c}{2}}-\beta)^{2}\cdot Var(\mathcal{L}_{A}).$ (26)
@@
According to the fourth assumption, we are interested in the range where $c\cdot(\beta+1)\geq 1$. Using the limits of $\beta$ defined in Eq. 9, we obtain $\beta\in\left[\frac{1-c}{c},\frac{2-c}{c}\right]$. As $\beta\to\frac{2-c}{c}$, the representation of the aligned samples is vanishing (according to Eq. 8) in the sampled batches, which leads to $\mathcal{L}_{A}>0$. Assuming a limited and non-vanishing variance $Var(\mathcal{L}_{A})$ (i.e., the model outputs non-constant predictions for samples from the majority group), we have:
@@
$\beta\to\frac{2-c}{c}\implies Var(\mathcal{L}_{sw})\approx 0,\text{ while }Var(\mathcal{L}_{ext})\neq 0$ for $c\in(0,1]$. This non-vanishing variance of GERNE’s loss, if controlled with tuning $\beta$ to ensure stability, gives the model the chance of escaping sharp minima similar to gradient random perturbation [1] and therefore, improves generalization [33, 23, 34].
Appendix C Simplifying the Bounds of $\beta$
We aim to simplify the upper and lower bounds of $\beta$ in Eq. 9. We start by simplifying the upper bound:
$\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}\neq\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2}),$
where $i_{ya}^{1}=-\frac{\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1,i_{ya}^{2}=\frac{1-\alpha_{ya}}{c\cdot(\frac{1}{A}-\alpha_{ya})}-1$, and under the following constraints: $\forall y\in\mathcal{Y},\sum_{a}\alpha_{ya}=1$, $\alpha_{ya}\in]0,1[\setminus\{\frac{1}{A}\}$, $A\geq 2$, and $c\in(0,1]$.
We note that $i_{ya}^{1}$ is a decreasing, and $i_{ya}^{2}$increasing function in $\alpha_{ya}$. We can show that if $\alpha_{ya}<\frac{1}{A}$, then $i_{ya}^{2}>i_{ya}^{1}$, and if $\alpha_{ya}>\frac{1}{A}$, then $i_{ya}^{1}>i_{ya}^{2}$. We conclude with the following:
$\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}<\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2})=i_{y^{\prime}a^{\prime}}^{2},\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\\ \alpha_{ya}>\frac{1}{A}\end{subarray}}\max(i_{ya}^{1},i_{ya}^{2})=i_{y^{\prime\prime}a^{\prime\prime}}^{1},$
where:
$\alpha_{y^{\prime}a^{\prime}}=\min_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\end{subarray}}\alpha_{ya}<\frac{1}{A},\alpha_{y^{\prime\prime}a^{\prime\prime}}=\max_{\begin{subarray}{c}{(y,a)\in\mathcal{G}}\end{subarray}}\alpha_{ya}>\frac{1}{A}.$
Since $\sum_{k}\alpha_{y^{\prime}k}=1$, we have:
$\sum_{k\neq a^{\prime}}\alpha_{y^{\prime}k}=1-\alpha_{y^{\prime}a^{\prime}},$
which implies that there exists $j\neq a^{\prime}$ such that:
$\alpha_{y^{\prime}j}\geq\frac{1-\alpha_{y^{\prime}a^{\prime}}}{A-1}.$
Given that $\alpha_{y^{\prime}a^{\prime}}<\frac{1}{A}$ and $A\geq 2$, it follows that
$\frac{1-\alpha_{y^{\prime}a^{\prime}}}{A-1}>\frac{1}{A},$
and hence $\alpha_{y^{\prime}j}>\frac{1}{A}$. Therefore, we have
$\max(i_{y^{\prime}j}^{1},i_{y^{\prime}j}^{2})=i_{y^{\prime}j}^{1}=-\frac{\alpha_{y^{\prime}j}}{c\cdot(\frac{1}{A}-\alpha_{y^{\prime}j})}-1.$
Since $i_{ya}^{1}$ is a decreasing function in $\alpha_{ya}\in(\frac{1}{A},1]$, we have
$i_{y^{\prime}j}^{1}\leq-\frac{\frac{1-\alpha_{y^{\prime}a^{\prime}}}{A-1}}{c\cdot(\frac{1}{A}-\frac{1-\alpha_{y^{\prime}a^{\prime}}}{A-1})}-1=\frac{1-\alpha_{y^{\prime}a^{\prime}}}{c\cdot(\frac{1}{A}-\alpha_{y^{\prime}a^{\prime}})}-1=i_{y^{\prime}a^{\prime}}^{2}.$
Thus: $i_{y^{\prime}j}^{1}\leq i_{y^{\prime}a^{\prime}}^{2}$, and since $i_{y^{\prime\prime}a^{\prime\prime}}^{1}\leq i_{y^{\prime}j}^{1}$, we conclude that the upper bound of $\beta$ is: $\beta_{\max}=i_{y^{\prime\prime}a^{\prime\prime}}^{1}$. For the lower bound of $\beta$, we can follow the same previous step, but we choose $\beta=-1$ as the lower bound ($\beta=-1$ satisfies Eq. 9 and simulates ERM training as shown in Sec. 4.1.3). In conclusion, for the known attributes case, we tune $\beta$ within the interval: $\beta_{\min},\beta_{\max}]=[-1,i_{y^{\prime\prime}a^{\prime\prime}}^{1}]$. The upper bound $\beta_{\max}$ is inversely proportional to both $c,A$. Consequently, larger values of $c$ reduce the feasible range for the extrapolation factor, making GERNE appear more sensitive to small variations in $\beta$.
Appendix D Algorithm 2
@@
Input: $\mathcal{X}_{y}\subseteq\mathcal{X}$ for $y\in\mathcal{Y}$, $f$ with initial $\theta=\theta_{0},\tilde{\theta}=\tilde{\theta_{0}}$ (parameters of the biased model $\tilde{f}$), #\# epochs $E$, batch size per class label $B$, #\# classes $K$, #\# attributes $\tilde{A}=2$, learning rate $\eta$.
1:Training $\tilde{f}$ on biased batches with class balanced accuracy CBA $\frac{1}{K}\sum_{y\in\mathcal{Y}}\mathbb{P}_{x|y}(y=\arg\max_{y^{\prime}\in\mathcal{Y}}\tilde{f}_{y^{\prime}}(x))\quad$ as the evaluation metric for model selection.
2:Select a threshold $t$ and create the pseudo-groups $\tilde{\mathcal{G}}$: For each class $y$, we compute the predictions $\tilde{y}_{i}=\text{softmax}(\tilde{f}(x_{i}))_{y}$ for each $x_{i}\in\mathcal{X}_{y}$. We then form the pseudo-minority group $\mathcal{X}_{y,\tilde{a}=1}$ as the samples $x_{i}$ that have the smallest $\lfloor t\cdot|\mathcal{X}_{y}|\rfloor$ predictions. The remaining samples form the pseudo-majority group $\mathcal{X}_{y,\tilde{a}=2}=\mathcal{X}_{y}\setminus\mathcal{X}_{y,\tilde{a}=1}$.
3:Follow Algorithm 1 with $\tilde{\mathcal{G}}$ replacing $\mathcal{G}$. We consider a higher upper bound for $\beta$ than $\beta_{\max}$ derived in Appendix C as justified by the proof of Proposition 1. in Appendix E.
@@
Creating both biased and less biased batches using the pseudo-groups $\mathcal{\tilde{G}}$, and with $\beta$ as a hyperparameter, we can simulate batches with a more controllable conditional attribute distribution. Specifically, for $(y,a)\in\mathcal{G}$, we can achieve scenarios where $p_{ext}(a|y)>\max_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)\text{ or }p_{ext}(a|y)<\min_{\tilde{a}\in\mathcal{\tilde{A}}}p(a|\tilde{a},y)$ as opposed to Eq. 13.
Proof. We define $\alpha_{y\tilde{a}}$ the same way as in Eq. 2 for the created pseudo-groups: $\alpha_{y\tilde{a}}=\frac{|\mathcal{X}_{y,\tilde{a}}|}{|\mathcal{X}_{y}|}$. For a constant $c$ and $A=2$, we create the less biased batch as in Eq. 4:
[…]
@@
[…]
@@
$p_{ext}(a|y)=\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\alpha_{y\tilde{a}}\cdot p(a|\tilde{a},y)+c\cdot(\beta+1)\cdot(\frac{1}{2}-\alpha_{y\tilde{a}})\cdot p(a|\tilde{a},y).$ (30)
For $p(a|\tilde{a}=1,y)\neq p(a|\tilde{a}=2,y)$ and $\alpha_{y1}\neq\frac{1}{2}$, to make $p_{ext}(a|y)=p$ for some $p\in[0,1]$, we tune $\beta$ until reaching the $\beta_{target}$ defined as follows:
$\beta_{target}=\frac{p-\sum_{\tilde{a}\in\tilde{\mathcal{A}}}\alpha_{y\tilde{a}}\cdot p(a|\tilde{a},y)}{\sum_{\tilde{a}\in\tilde{\mathcal{A}}}c\cdot(\frac{1}{2}-\alpha_{y\tilde{a}})\cdot p(a|\tilde{a},y)}-1.$ (31)
Discussion. When $\alpha_{y1}=\frac{1}{2}$, our algorithm is equivalent to sampling uniformly from $\mathcal{X}_{y}$ and equally from classes. When $p(a|\tilde{a}=1,y)=p(a|\tilde{a}=2,y)$, it implies that $\tilde{f}$ has distributed the samples with attribute $a$ and class $y$ equally between the two pseudo-groups $\mathcal{X}_{y,\tilde{a}\in\mathcal{A}}$. However, in practice, this is exactly the scenario that $\tilde{f}$ is designed to avoid. Specifically, if $a$ represents the presence of spurious attributes (i.e., the majority group), it is likely that $p(a|\tilde{a}=1,y)<p(a|\tilde{a}=2,y)$. Conversely, when $a$ represents the absence of spurious features (i.e., the minority group), we would expect $p(a|\tilde{a}=1,y)>p(a|\tilde{a}=2,y)$. In fact, $\tilde{f}$ is explicitly trained to exhibit a degree of bias, which inherently disrupts the above equality.